Ticket Number: HEA-NOC/19980702-02 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> MCI
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19980702 13:30 UTC+1
Ticket Closed: 19980708 10:30:32 UTC Problem End: 19980708 09:40 UTC+1
The three E1 circuits connecting HEAnet to MCI in the U.S. were
unreliable. MCI is continuing to test.
Traffic to and from the US was impeded. In particular, short(10-20sec)
flaps on the line caused long (<=30 min) dips in access to the rest of
the internet, caused by route dampening.
Reported to MCI: fault #943 02Jul98.
Update 09:00 03Jul98: Full service came back about 17:00 yesterday,
but no update from MCI. We've been hit by emergency maintenance
in West Orange so we're back down to 2MB for the moment.
Update 17:45 03Jul98: MCI continuing to test. Operating at 4MB since
13:30. An error on MCI's part caused an outage at 17:30, which
was resolved at 17:39 by the NOC.
19980705 13:40 MCI restore mapping to E1-1, we're now running
at 6Mbps. Will review position with MCI on 6th July.
19980708 11:30: MCI replaced card in West Orange bordercore router,
ending the flaps. Each of the E1 lines were tested separately,
and no faults were found; ticket has been closed.
Time to Fix:
First report to resolution: 5d 20h 10m.